# Conditional and Modal Reasoning in Large Language Models

### v3

@@ Abstract
The reasoning abilities of large language models (LLMs) are the topic of a growing body of research in AI and cognitive science. In this paper, we probe the extent to which twenty-five LLMs are able to distinguish logically correct inferences from logically fallacious ones. We focus on inference patterns involving conditionals (e.g., ‘If Ann has a queen, then Bob has a jack’) and epistemic modals (e.g., ‘Ann might have an ace’, ‘Bob must have a king’). These inferences have been of special interest to logicians, philosophers, and linguists, since they play a central role in the fundamental human ability to reason about distal possibilities. Assessing LLMs on these inferences is thus highly relevant to the question of how much the reasoning abilities of LLMs match those of humans. Among the LLMs we tested, all but the GPT-4 model family often make basic mistakes with conditionals, though zero-shot chain-of-thought prompting helps them make fewer mistakes. Moreover, even the GPT-4 family displays logically inconsistent judgments across inference patterns involving epistemic modals, and almost all models give answers to certain complex conditional inferences widely discussed in the literature that do not match human judgments. These results highlight gaps in basic logical reasoning in today’s LLMs.
@@ 1 Introduction
Figure 1: Summary of performance on the simple logical inference patterns discussed in § 4. Guessing accuracy is at 50%. Larger models generally perform better, and most models show clear weakness at this task.
@@
After providing some background in § 2 and detailing our experimental setup in § 3, we discuss results for a number of inference patterns in § 4. We find that among the LLMs tested, all but the GPT-4 family often make basic mistakes with conditionals. Moreover, even GPT-4 displays logically inconsistent judgments across inference patterns involving epistemic modals, and almost all models give answers to certain complex conditional inferences that do not match reported human judgments. We also show that models’ performance on our reasoning tasks is highly correlated with that of Chatbot Arena Elo ratings Chiang et al. (2024), MMLU Hendrycks et al. (2020), and GSM8K Cobbe et al. (2021), supporting the hypothesis that logical reasoning abilities are predictive of general model capabilities and performance on downstream tasks. In sum, our main contributions in this paper are:
@@ 2.1 Logical inference
Regarding purely logical reasoning, a series of benchmarks have been created in recent years Tafjord et al. (2020); Tian et al. (2021); Han et al. (2022); Saparov and He (2023); Saparov et al. (2023), and various strategies have been proposed to solve some of them Creswell et al. (2023); Kazemi et al. (2023); Olausson et al. (2023); Pan et al. (2023); Poesia et al. (2023); Ye et al. (2023). Those benchmarks primarily focus on multi-step reasoning, where a proof is required from premises to the hypothesis. Here we target single-step inference patterns, which we treat as more fundamental. The inability to recognize the basic inference patterns we study here could provide further explanations of failures on those multi-step reasoning problems. Additionally, none of the work above studies modal operators and their interactions with conditionals.
@@ 2.2 Modals and conditionals
Here we draw specifically on the logical and philosophical literature on modals and conditionals. Enormous progress has been made in the last half century on both topics. First, modal operators like ‘must’ and ‘might’ have been successfully modeled as quantifiers over possible worlds Kripke (1963); Kratzer (1981). That is, just as ‘Every boy is sitting’ quantifies universally over all boys (in a given domain), ‘It must be raining’ quantifies over all possible worlds (in a given domain) and says that it is raining in all of them; and just as ‘Some boy is sitting’ quantifies existentially over boys, ‘It might be raining’ says that it is raining in some possible world. This interpretation yields corresponding logics of modality, with the details depending on how the domain of possible worlds is obtained (and on the interpretation of the other connectives and operators with which modals interact).
@@
These points are worth emphasizing, since although the material analysis is almost universally rejected by theorists of the conditional, it is still assumed in much existing work testing the logical capacities of humans and LLMs, in both cognitive science and artificial intelligence (e.g., in the recent Wan et al. (2024), which treats the material analysis as one of the benchmarks of correct reasoning with conditionals). This is a serious blindspot, since failing to reason in accord with the material conditional may be logically correct; and, conversely, reasoning in accord with the material conditional may be a serious logical mistake.
The most popular alternative treats ‘if $p$, then $q$’ as a restricted modal operator, which says that $q$ is true in all $p$-worlds (in a given domain). As for modals, this yields corresponding logics, with the details again depending on assumptions about which $p$-worlds are in the domain, together with the interpretation of other connectives Stalnaker (1968); Lewis (1973b); Egré and Rott (2021).
Although the material analysis is almost universally rejected, there is ongoing controversy about the correct logic of conditionals and modals. We have chosen a wide range of inference patterns to test: in many of these cases there is (near) universal agreement about whether the inference pattern is valid. In other cases, there is less agreement about whether the pattern is truly valid, but even in those cases, there is for the most part agreement about whether naive human reasoners are inclined to draw the inference, and the remaining controversy is about how to model those patterns (as genuine (in)validities or the result of systematic context-shifting). We do not aim to take a position in these complex debates here but rather to compare the behavior of LLMs to widely reported human inferential dispositions. In future work, we plan to compare the behavior of LLMs with human subjects (compare the methodology of Pavlick and Kwiatkowski, 2019; Dasgupta et al., 2022; Webson et al., 2023); in this paper, we compare LLMs against expert claims about inference from the philosophical/logical literature.
@@ 2.3 Natural language inference
Valid Inferences Examples Valid Modal Inferences Examples
8pt. DS: Either Fido is inside or Fido is in the garden MiN: Mary might not have been at the wedding
$p\vee q,\neg q\vdash p$ Fido is not in the garden $\lozenge\neg p\vdash\neg\Box p$ $\vdash$ It’s not the case that Mary must have been at the wedding
$\vdash$ Fido is inside
MP: If Mary was at the wedding, then Sue was at the wedding. NMu: It’s not the case that Mary must have been at the wedding
$p\to q,p\vdash q$ Mary was at the wedding $\neg\Box p\vdash\lozenge\neg p$ $\vdash$ Mary might have been at the wedding
$\vdash$ Sue was at the wedding.
MT: If Mary was at the wedding, then Sue was at the wedding.
$p\to q,\neg q\vdash\neg p$ Sue was not at the wedding.
Mary was not at the wedding
8pt. Invalid Inferences Controversial Modal Inferences
@@
$p\to q,q\vdash p$ Sue was at the wedding $p\vee\Box q,\neg\Box q\vdash\neg p$ It’s not the case that Fido must be in the garden
$\vdash$ Mary was at the wedding $\vdash$ Fido isn’t inside
CONV: If Mary was at the wedding, then Sue was at the wedding. DSmi: Either Fido is inside or Fido must be in the garden
$p\to q\vdash q\to p$ $\vdash$ If Sue was at the wedding, then Mary was at the wedding $p\vee\Box q,\lozenge\neg q\vdash\neg p$ Fido might not be in the garden.
$\vdash$ Fido isn’t inside.
DA: If Mary was at the wedding, then Sue was at the wedding MTmi: If Mary was at the wedding, then Sue must have been there.
@@
$\vdash$ Sue was not at the wedding $\vdash$ Mary was not at the wedding.
@@
8pt. Controversial Inferences
8pt. AS: If the match is struck, then it will light. CMP: If the warriors don’t win, then
$p\to q\vdash(p\wedge r)\to q$ $\vdash$ If the match is struck and has been soaked in water, then it will light. $p\to(q\to r),p$ if the Lakers don’t win, the Celtics will.
CT: If it’s raining, then it’s not raining hard. $\vdash q\to r$ The Warriors won’t win.
$p\to q\vdash\neg q\to\neg p$ $\vdash$ If it’s raining hard, then it’s not raining. $\vdash$ If the Lakers don’t win, the Celtics will.
@@ 3.1 Models
We tested the logical inference judgments of the 25 LLMs listed in Figure 1, including both open and closed ones, cited in Appendix A. The Anthropic, Google, and OpenAI models were run through their respective APIs. For the local models, we used LM Studio to create local inference servers on an Apple M2 Ultra Mac Studio. All local models were 6-bit quantized versions in the Q6_K GGUF format available on Hugging Face. All experiments took $\sim$250 GPU hours and $\sim$\$1,000 for API calls. A Jupyter notebook and all data is available at github.com/wesholliday/llm-logic.
@@ 3.2 Data
For our experiments, we created a bank of questions for dozens of inference patterns, allowing us to individually probe LLMs’ inferential capacities with respect to each inference pattern. For each of these, we began by handcrafting a paradigm instance. E.g., for Modus Tollens (MT) (see subsection 2.3), our paradigm instance was: “From ‘If Mary was at the wedding, then Sue was at the wedding’ together with ‘Sue was not at the wedding’, can we infer ‘Mary was not at the wedding’?” With Claude 2, which was not one of the LLMs we evaluated, we created 19 additional instances of the pattern, e.g., “From ‘If the alien visited Mars, then the robot visited Jupiter’ together with ‘The robot did not visit Jupiter’, can we infer ‘The alien did not visit Mars’?” To guard against an LLM judging an inference based on world knowledge rather than the inference’s logical form, we also used Claude 2 to create 20 instances with nonsense predicates, e.g., “From ‘If the flugel was blimmed, then the flugel was zargled’ together with ‘The flugel was not zargled’, can we infer ‘The flugel was not blimmed’?” We denote the version of an inference with nonsense predicates with an ‘x’ at the end of its name, e.g., MTx. We also tested order effects for all inferences with two premises by switching the order of premises in an ‘o’ version. Using nonsense predicates and switched premise order yields an ‘ox’ version that we also tested.
@@ 3.3 Evaluation
For each of the 20 instances of an inference pattern question and each LLM, we posed the instance to the LLM with temperature 0 and then 1, along with a prompt for either zero-shot, few-shot, or zero-shot chain-of-thought Kojima et al. (2022) conditions (see Appendix B). For temperature $=1$, we asked the question repeatedly to get an empirical frequency for ‘yes’ and for ‘no’. [^1] We used the following stopping rule: if for a particular question the LLM answered ‘yes’ 10 times in a row or ‘no’ 10 times in a row, then we proceeded to the next question; otherwise we posed the same question a total of 20 times to the model. For GPT-4, Gemini 1.5 Pro, and Claude 3 Opus, we reduced 10 to 5 to reduce cost. For OpenAI models, the empirical frequencies obtained closely agree with the log probabilities, but log probs were not available for all models.
A natural question is whether 20 instances of each inference pattern is enough. To address this, we looked at the Pearson correlation coefficients between the yes-frequency of model responses to the 20 instances of an inference with nonsensical predicates (like MTx) and the yes-frequency of model responses to the 20 instances with sensical predicates (like MT). The overall correlation for all inferences in Table 2.3 is .85. While we could run more instances for all inference patterns, limited only by time and cost, the high correlations between responses to sensical and nonsensical instances suggest that we already have enough instances to see how an LLM responds to the logical form in question, not just to contingent features of particular instances of that form.
We also tested to what extent our results are sensitive to the choice of the word ‘infer’ in our prompts, as opposed to other phrases we consider equivalent in this context: ‘deduce’, ‘conclude’, ‘logically infer’, ‘logically deduce’, and ‘logically conclude’. Could it be, e.g., that models might commit fewer conditional fallacies if we prompt them with one of these other phrases? We ran the instances of the AC inference (see subsection 2.3) using each of the mentioned substitutes for ‘infer’. The results are qualitatively the same as for ‘infer’, and the correlation coefficients between the yes-frequencies for ‘infer’ and for each of the substitutes are over .9.9.9.9. Thus, which of the phrases above we use apparently makes little difference. [^2] By contrast, there was a bigger difference when we explicitly asked LLMs, “Is this form of argument logically valid?” (we call this the ‘v’ variant of each question). This decreased the rate at which some LLMs accepted fallacious inferences, perhaps due to the presence of texts on formal logic in the training data. However, it also decreased the rate at which some LLMs accepted valid inferences (see Supplementary Material). In any case, the qualitative observations we make still apply with this alternative formulation of the questions.
@@ 4 Results
The inferences we tested are summarized in subsection 2.3. To establish a baseline for performance, we tested a variety of inferences whose (in)validity is uncontroversial (as labeled in the table): MP, MT, AC, DA, DS, and x, o, and ox variants thereof, as well as INV, CONV, and x variants thereof. The aggregate performance of our models on these inferences is reported in Table 2 for the temperature 0 settings, in addition to Figure 1 at the beginning of the paper. We see that the results significantly vary across models, and performance roughly correlates with model size. No open-weight model achieves above 80% accuracy on this set of tasks. Among closed-source models, the GPT-4 models outperform Claude 3 Opus and Gemini 1.5 Pro.
@@
GPT-4 Turbo (2024-04-09) 98.5 +0.2 +0.6
GPT-4 Turbo (1106) 96.5 +0.0 +2.1
GPT-4 (0613) 95.6 +0.2 +3.1
Gemini 1.5 Pro 92.1 -8.3 +6.7
GPT-4 (0314) 91.7 +5.8 +7.7
GPT-4o (2024-05-13) 87.9 +1.3 +9.0
Claude 3 Opus 82.9 +6.9 +14.6
Gemini 1.5 Flash 82.1 -3.3 +8.8
Llama 3 Instruct 70B 77.1 +5.4 +13.1
Mixtral 8x7B 75.8 -0.4 +12.9
Claude 3 Sonnet 74.6 +4.0 +10.2
Llama 3 Instruct 8B 69.8 -3.1 +11.3
Claude 3 Haiku 69.0 -3.1 +7.7
Phi-2 60.0 -7.7 -8.1
Mistral 7B 59.8 -4.6 +9.2
Code Llama 13B 59.0 -9.0 +7.7
Code Llama 7B 57.5 +5.6 +0.2
GPT-3.5 Turbo (0613) 56.3 +1.7 +13.8
Code Llama 34B 56.3 -6.5 +4.4
GPT-3.5 Turbo (0125) 56.0 +2.1 +22.5
Llama 2 Chat 13B 55.2 +2.1 +0.0
Llama 2 Chat 70B 55.2 -3.1 +3.8
Llama 2 Chat 7B 55.0 -5.2 -3.5
GPT-3.5 Turbo (1106) 54.0 -0.4 +15.8
Yi Chat 34B 45.6 +4.4 +20.4
Table 2: Model performance on some simple inferences with different prompting setups.
We observe that the models perform similarly in the few-shot setting compared to the zero-shot setting (paired t-test not statistically significant). This is presumably because the few-shot examples we use (see Appendix B) are about conjunctions and disjunctions, not directly about conditionals and modals. The motivation of the few-shot setting is to make the logical reasoning task clear in context. The results suggest that models understand the task requirement in the zero-shot setting, yet they are not capable of consistently recognizing valid and invalid inference patterns. On the other hand, zero-shot chain-of-thought (CoT) prompting does dramatically improve performance (paired t-test statistically significant). The GPT-4 models, Gemini 1.5 Pro, and Claude 3 Opus all achieve near perfect accuracy (98.6%+), and Llama 3 70B is able to achieve 90%+ accuracy, now comparable to Gemini 1.5 Flash. These results contribute to the cumulative evidence that CoT elicits and improves reasoning in LLMs Wei et al. (2022); Kojima et al. (2022); Suzgun et al. (2022). The performance trends of the temperature 1 settings are similar, and we include figures for those in Appendix C. Next, we summarize some noteworthy findings based on specific inferences from Table 2.3.

### v4

@@ Abstract
The reasoning abilities of large language models (LLMs) are the topic of a growing body of research in AI and cognitive science. In this paper, we probe the extent to which twenty-nine LLMs are able to distinguish logically correct inferences from logically fallacious ones. We focus on inference patterns involving conditionals (e.g., ‘If Ann has a queen, then Bob has a jack’) and epistemic modals (e.g., ‘Ann might have an ace’, ‘Bob must have a king’). These inferences have been of special interest to logicians, philosophers, and linguists, since they play a central role in the fundamental human ability to reason about distal possibilities. Assessing LLMs on these inferences is thus highly relevant to the question of how much the reasoning abilities of LLMs match those of humans. All the LLMs we tested make some basic mistakes with conditionals or modals, though zero-shot chain-of-thought prompting helps them make fewer mistakes. Even the best performing LLMs make basic errors in modal reasoning, display logically inconsistent judgments across inference patterns involving epistemic modals and conditionals, and give answers about complex conditional inferences that do not match reported human judgments. These results highlight gaps in basic logical reasoning in today’s LLMs.
@@ 1 Introduction
Figure 1: Summary of performance on the uncontroversial logical inference patterns discussed in § 4. Guessing accuracy is 50%. Larger models generally perform better, and most models show clear weakness at this task.
@@
After providing some background in § 2 and detailing our experimental setup in § 3, we discuss results for a number of inference patterns in § 4. We find that among the LLMs tested, all models commit basic fallacies in reasoning with modals and conditionals. Even the best performing models display logically inconsistent judgments across certain inference patterns involving modals and conditionals. And almost all models give answers to certain complex conditional inferences that do not match reported human judgments. We also show that models’ performance on our reasoning tasks is highly correlated with that of Chatbot Arena Elo ratings Chiang et al. (2024), MMLU Hendrycks et al. (2020), and GSM8K Cobbe et al. (2021), supporting the hypothesis that logical reasoning abilities are predictive of general model capabilities and performance on downstream tasks. In sum, our main contributions in this paper are:
@@ 2.1 Logical inference
Regarding purely logical reasoning, a series of benchmarks have been created in recent years Tafjord et al. (2020); Tian et al. (2021); Han et al. (2022); Saparov and He (2023); Saparov et al. (2023), and various strategies have been proposed to solve some of them Creswell et al. (2023); Kazemi et al. (2023); Olausson et al. (2023); Pan et al. (2023); Poesia et al. (2023); Ye et al. (2023). Those studies (with the exceptions of Tafjord et al. 2020; Creswell et al. 2023) primarily focus on multi-step reasoning, where a proof is required from premises to the hypothesis. Here we target single-step inference patterns, which we treat as more fundamental. Indeed, the inability to recognize the basic inference patterns we study here could provide further explanations of failures on multi-step reasoning problems. Most importantly, none of the work above studies modal operators, conditional operators, and their interactions, our novel focus in this paper.
@@ 2.2 Modals and conditionals
Here we draw specifically on the logical and philosophical literature on modals and conditionals. Enormous progress has been made in the last half century on both topics. First, modal operators (like ‘must’ and ‘might’) have been successfully modeled as quantifiers over possible worlds Kripke (1963); Kratzer (1981). That is, just as ‘Every boy is sitting’ quantifies universally over all boys (in a given domain), ‘It must be raining’ quantifies over all possible worlds (in a given domain—in this case, an epistemic domain) and says that it is raining in all of them; and just as ‘Some boy is sitting’ quantifies existentially over boys, ‘It might be raining’ says that it is raining in some epistemically possible world. Similarly, a deontic modal like ‘may’ can be interpreted as quantifying over a domain of deontically possible worlds, so that ‘You may eat a cookie’ is true just in case you eat a cookie in some deontically possible world, i.e., one where all the actual deontic requirements are satisfied. Likewise, ‘must’, on its deontic interpretation (as in ‘You must eat this cookie’) quantifies universally over deontically possible worlds, and says that you eat cookies in all of them.
This interpretation of modals yields corresponding logics of modal language, with the details depending on how the domain of possible worlds is obtained (and the interpretation of the other connectives and operators with which modals interact).
@@
These points are worth emphasizing, since although the material analysis is almost universally rejected by theorists of the conditional, it is still assumed in much existing work testing the logical capacities of humans and LLMs, in both cognitive science and artificial intelligence (e.g., in the recent Wan et al. 2024, which treats the material analysis as one of the benchmarks of correct reasoning with conditionals). This is a serious blindspot, since failing to reason in accord with the material conditional may be logically correct; and, conversely, reasoning in accord with the material conditional may be a serious logical mistake.
The most popular alternative treats ‘if $p$, then $q$’ as a restricted modal operator, which says that $q$ is true in all $p$-worlds (in a given domain). Just as for modals, this yields corresponding logics, with the details again depending on assumptions about which $p$-worlds are in the domain, together with the interpretation of other connectives Stalnaker (1968); Lewis (1973b); Egré and Rott (2021).
Although the material analysis is almost universally rejected, there is ongoing controversy about the correct logic of conditionals and modals. We have chosen a wide range of inference patterns to test: in many of these cases there is (near) universal agreement about whether the inference pattern is valid. In other cases, there is less agreement about whether the pattern is truly valid, but even in those cases, there is for the most part agreement about whether typical human reasoners are inclined to draw the inference, and the remaining controversy is about how to model those patterns (as genuine (in)validities or the result of systematic shifts in interpretation). We do not aim to take a position in these complex debates here but rather to compare the behavior of LLMs to widely reported human inferential dispositions. In future work, we plan to compare the behavior of LLMs with human subjects (compare the methodology of Pavlick and Kwiatkowski, 2019; Dasgupta et al., 2022; Webson et al., 2023); here we compare LLMs against expert claims about inference from the literature in philosophy, logic, and semantics.
@@ 2.3 Natural language inference
Valid Inferences Examples Controv. Conditional Examples
8pt. DS: Either Fido is inside or Fido is in the garden. Fido is not in the garden. CT: If it’s raining, then it’s not raining hard.
$p\vee q,\neg q\vdash p$ $\vdash$ Fido is inside. $p\to q\vdash\neg q\to\neg p$ $\vdash$ If it’s raining hard, then it’s not raining.
MP: If Mary was at the wedding, then Sue was at the wedding. AS: If the match is struck, then it will light.
$p\to q,p\vdash q$ Mary was at the wedding. $p\to q\vdash(p\wedge r)\to q$ $\vdash$ If the match is struck and has been soaked in water,
$\vdash$ Sue was at the wedding. then it will light.
MT: If Mary was at the wedding, then Sue was at the wedding. CMP: If the Warriors don’t win, then
$p\to q,\neg q\vdash\neg p$ Sue was not at the wedding. $\vdash$ Mary was not at the wedding. $p\to(q\to r),p$ if the Lakers don’t win, the Celtics will.
MiN: Mary might not have been at the wedding. $\vdash q\to r$ The Warriors won’t win.
$\lozenge\neg p\vdash\neg\Box p$ $\vdash$ It’s not the case that Mary must have been at the wedding. $\vdash$ If the Lakers don’t win, the Celtics will.
NMu: It’s not the case that Mary must have been at the wedding.
$\neg\Box p\vdash\lozenge\neg p$ $\vdash$ Mary might not have been at the wedding.
8pt. Invalid Inferences Controv. Modal
@@
$p\to q,q\vdash p$ Sue was at the wedding. $p\vee\Box q,\neg\Box q\vdash p$ It’s not the case that Fido must be in the garden.
$\vdash$ Mary was at the wedding $\vdash$ Fido is inside.
CONV: If Mary was at the wedding, then Sue was at the wedding. DSmi: Either Fido is inside or Fido must be in the garden.
$p\to q\vdash q\to p$ $\vdash$ If Sue was at the wedding, then Mary was at the wedding. $p\vee\Box q,\lozenge\neg q\vdash p$ Fido might not be in the garden.
$\vdash$ Fido is inside.
DA: If Mary was at the wedding, then Sue was at the wedding. MTmi: If Mary was at the wedding, then Sue must have been there.
@@
$\vdash$ Sue was not at the wedding. $\vdash$ Mary was not at the wedding.
@@
MuDistOr: The envelope must have been upstairs or under a bed. WSFC: John might clean his room or he might go to the lecture.
$\Box(p\!\vee\!q)\!\vdash\!\Box p\!\vee\!\Box q$ $\vdash$ The envelope must have been upstairs or it must have been under a bed. $\lozenge p\vee\lozenge q\vdash\lozenge p\wedge\lozenge q$ $\vdash$ John might clean his room and he might go to the lecture.
MiAg: The envelope might be upstairs and the envelope might be under a bed. NSFC: John might clean his room or go to the lecture.
$\lozenge p\!\wedge\!\lozenge q\vdash\!\lozenge(p\!\wedge\!q)$ $\vdash$ The envelope might be upstairs under a bed. $\lozenge(p\vee q)\vdash\lozenge p\wedge\lozenge q$ $\vdash$ John might clean his room and John might go to the lecture.
@@ 3.1 Models
We tested the logical inference judgments of the 29 LLMs listed in Figure 1, including both open and closed ones, cited in Appendix A. The Anthropic, Google, Mistral, and OpenAI models were run through their respective APIs. Open-weight models were run through cloud providers (together.ai and fireworks.ai). All experiments cost $\sim$\$3,000 for API calls. All code and data for the experiments are available at github.com/wesholliday/llm-logic, which also includes data for OpenAI’s o1 models (released after this paper was submitted).
@@ 3.2 Data
For our experiments, we created a bank of questions for dozens of inference patterns, allowing us to individually probe LLMs’ inferential capacities with respect to each inference pattern. For each of these, we began by handcrafting a paradigm instance. E.g., for Modus Tollens (MT) (see subsection 2.3), our paradigm instance was: “From ‘If Mary was at the wedding, then Sue was at the wedding’ together with ‘Sue was not at the wedding’, can we infer ‘Mary was not at the wedding’?” With the help other other LLMs (mostly Claude 2, supplemented by GitHub Copilot and Devin), we created 19 additional instances of the pattern, e.g., “From ‘If the alien visited Mars, then the robot visited Jupiter’ together with ‘The robot did not visit Jupiter’, can we infer ‘The alien did not visit Mars’?” To guard against an LLM judging an inference based on world knowledge rather than the inference’s logical form, we also created 20 instances with nonsense predicates, e.g., “From ‘If the flugel was blimmed, then the flugel was zargled’ together with ‘The flugel was not zargled’, can we infer ‘The flugel was not blimmed’?” We denote the version of an inference with nonsense predicates with an ‘x’ at the end of its name, e.g., MTx. We also tested order effects for all inferences with two premises by switching the order of premises in an ‘o’ version. Using nonsense predicates and switched premise order yields an ‘ox’ version that we also tested. We reviewed all LLM-generated stimuli and made necessary adjustments by hand.
@@ 3.3 Evaluation
For each of the 20 instances of an inference pattern question and each LLM, we posed the instance to the LLM with temperature 0 and then 1, along with a prompt for either zero-shot, few-shot, or zero-shot chain-of-thought Kojima et al. (2022) conditions (see Appendix B). For temperature $=1$, we asked the question repeatedly to get an empirical frequency for ‘yes’ and for ‘no’. [^1] We used the following stopping rule: if for a particular question the LLM answered ‘yes’ 10 times in a row or ‘no’ 10 times in a row, then we proceeded to the next question; otherwise we posed the same question a total of 20 times to the model. For the models with the higher costs per token, we reduced 10 to 5 to reduce cost. For OpenAI models, the empirical frequencies obtained closely agree with the log probabilities, but log probs were not available for all models.
To assess the sufficiency of using 20 instances of each inference pattern, we looked at the Pearson correlation coefficients between the yes-frequency of model responses to the 20 instances of an inference with nonsensical predicates (like MTx) and the yes-frequency of model responses to the 20 instances with sensical predicates (like MT). The overall correlation for all inferences in Table 2.3 is .85. While we could run more instances for all inference patterns, limited only by time and cost, the high correlations between responses to sensical and nonsensical instances suggest that we already have enough instances to see how an LLM responds to the logical form in question, not just to contingent features of particular instances of that form.
We also tested to what extent our results are sensitive to the choice of the word ‘infer’ in our prompts, as opposed to other phrases we consider equivalent in this context: ‘deduce’, ‘conclude’, ‘logically infer’, ‘logically deduce’, and ‘logically conclude’. Could it be, e.g., that models might commit fewer conditional fallacies if we prompt them with one of these other phrases? We ran the instances of the AC inference (see subsection 2.3) using each of the mentioned substitutes for ‘infer’. The results are qualitatively the same as for ‘infer’, and the correlation coefficients between the yes-frequencies for ‘infer’ and for each of the substitutes are over .9.9.9.9. Thus, which of the phrases above we use apparently makes little difference. [^2] By contrast, there was a bigger difference when we explicitly asked LLMs, “Is this form of argument logically valid?” (we call this the ‘v’ variant of each question). This decreased the rate at which some LLMs accepted fallacious inferences, perhaps due to the presence of texts on formal logic in the training data. However, it also decreased the rate at which some LLMs accepted valid inferences (see the GitHub repository). In any case, the qualitative observations we make still apply with this alternative formulation of the questions.
@@ 4 Results
Key inferences we tested are summarized in subsection 2.3. To establish a baseline for performance, we tested a variety of inferences whose (in)validity is uncontroversial (as labeled in the left-hand side of the table): DS, MP, MT, AC, DA, and x, o, and ox variants thereof, as well as INV, CONV, MiN, NMu, MuDistOr, MiAg, and x variants thereof. The aggregate performance of our models on these inferences is reported in Table 2 for the temperature 0 settings, in addition to Figure 1 at the beginning of the paper. We see that the results significantly vary across models, and performance roughly correlates with model size. No model achieves above 90% accuracy on this set of tasks. Moreover, Llama 3.1 405B and 70B are the only open-weights models that achieve 80%+ accuracy, with the former being the best performer overall. The GPT-4 model family show generally strong performance.
@@
Llama 3.1 Instruct 405B 89.50 0.17 1.00
GPT-4 Turbo (2024-04-09) 87.17 2.00 4.33
Claude 3.5 Sonnet 85.33 2.83 2.33
GPT-4 Turbo (1106) 84.67 1.33 1.67
GPT-4 (0613) 83.50 2.50 5.17
Llama 3.1 Instruct 70B 81.33 3.33 5.50
Gemini 1.5 Pro 81.17 -4.83 7.67
GPT-4 (0314) 80.33 5.33 7.50
GPT-4o (2024-05-13) 79.67 0.50 9.00
GPT-4o mini 75.33 8.17 13.50
Claude 3 Opus 74.00 6.00 15.83
Gemini 1.5 Flash 72.83 -2.33 13.50
Mistral Large 2 72.17 10.67 12.50
Mixtral 8x7B 71.17 -7.83 0.17
Llama 3 Instruct 70B 70.33 5.33 11.00
Claude 3 Sonnet 68.17 2.17 12.00
Claude 3 Haiku 65.50 -3.67 11.00
Mixtral 8x22B 65.17 12.50 21.50
Gemma 2 27B 64.67 1.33 16.50
Llama 3 Instruct 8B 63.83 -4.00 8.17
Code Llama 34B 58.67 -7.83 2.67
GPT-3.5 Turbo (0125) 57.33 2.33 17.00
Llama 2 Chat 7B 55.00 -1.00 3.50
GPT-3.5 Turbo (1106) 54.33 1.17 14.67
Llama 3.1 Instruct 8B 53.67 8.00 19.50
Code Llama 7B 50.33 4.50 12.67
Llama 2 Chat 13B 49.67 7.83 9.00
Mistral 7B 49.17 0.17 21.17
Code Llama 13B 45.33 -8.00 8.17
Table 2: Model performance on uncontroversial inferences with different prompting setups.
We observe that the models perform similarly in the few-shot setting compared to the zero-shot setting (paired t-test not statistically significant). This is presumably because the few-shot examples we use (see Appendix B) are about conjunctions and disjunctions, not directly about conditionals and modals. The motivation of the few-shot setting is to make the logical reasoning task clear in context. The results suggest that models understand the task in the zero-shot setting, yet they are not capable of consistently recognizing valid and invalid inference patterns. On the other hand, zero-shot chain-of-thought (CoT) prompting does dramatically improve performance (paired t-test statistically significant): the best models all achieve accuracy near 90%, while still systematically making some mistakes with modal or conditionals. These results contribute to the cumulative evidence that CoT elicits and improves reasoning in LLMs Wei et al. (2022); Kojima et al. (2022); Suzgun et al. (2022), while pointing to lacunae that persist even with CoT. The performance trends of the temperature 1 settings are similar, and we include figures for those, and for CoT prompting, in Appendix C. Next, we summarize some noteworthy findings based on specific inferences from Table 2.3.